# Tropical Gradient Descent

### v1

@@ Abstract
We propose a gradient descent method for solving optimisation problems arising in settings of tropical geometry—a variant of algebraic geometry that has become increasingly studied in applications such as computational biology, economics, and computer science. Our approach takes advantage of the polyhedral and combinatorial structures arising in tropical geometry to propose a versatile approach for approximating local minima in tropical statistical optimisation problems—a rapidly growing body of work in recent years. Theoretical results establish global solvability for 1-sample problems and a convergence rate of $O(1/\sqrt{k})$. Numerical experiments demonstrate the method’s superior performance over classical descent for tropical optimisation problems which exhibit tropical convexity but not classical convexity. Notably, tropical descent seamlessly integrates into advanced optimisation methods, such as Adam, offering improved overall performance.
@@ 1 Introduction
Where algebraic geometry studies geometric properties that can be expressed algebraically of solution sets of systems of multivariate polynomials, tropical geometry restricts to the case of polynomials defined by linearising operations, where the sum of two elements is their maximum and the product of two elements is their sum. Evaluating polynomials with these operations results in piecewise linear functions. Tropical geometry is a relatively young field of pure mathematics established in the 1990s (with roots dating back further) and has recently become an area of active interest in computational and applied mathematics for its relevance to the problem of dynamic programming in computer science (Maclagan & Sturmfels, (2021)); in mechanism design for two-player games in game theory (Lin & Tran, (2019)); and statistical analyses of phylogenetic trees (Monod et al., (2022)). Given the connection of tropical geometry to applied sciences, developing and implementing computational techniques for tropical geometric settings is becoming increasingly important.
A central task to a vast majority of computational tasks and training machine learning models entails solving an optimisation problem. In this paper, we are motivated by the optimisation problems in statistics in tropical geometric settings. We focus on the classical gradient descent approach for optimisation and adapt it to the setting of tropical geometry. We thus provide a unifying gradient descent framework to solve the optimisation problems entailed in the computation of a wide range of statistical quantities and methods in tropical spaces which are an area of active interest in applied tropical geometry.
@@
The remainder of this paper is organised as follows. The following section contains the necessary foundations of tropical geometry to present, an introduction to tropical location problems functions (Comăneci, (2023)), and the array of statistical optimisation tasks studied in this paper. In Section 3 we formulate steepest descent with respect to the tropical norm and present the theoretical guarantees for tropical descent, most notably its convergence for a wide class of tropically quasi-convex functions and a convergence rate of $O(1/\sqrt{k})$. In Section 4, we perform a comprehensive numerical study of our proposed tropical gradient descent method and its performance in the statistical optimisation problems of interest. We conclude with a discussion of possible future work on tropical gradient methods in Section 5.
@@ 2 Background and Preliminaries
In this section, we overview the background on tropical geometry and present the statistical concepts and methodologies which are defined by an optimisation problem that we will study in the tropical setting.
@@ 2.1 Tropical Geometry
Here, we review the algebraic and geometric structure of the tropical projective torus, the state space for phylogenetic data and the domain of our tropical statistical optimisation problems.
@@
The $n$-dimensional tropical projective torus is a quotient space constructed by endowing $\mathbb{R}^{n+1}$ with the equivalence relation
@@
it is denoted by $\mathbb{R}^{n+1}/\mathbb{R}\mathbf{1}$. The generalised Hilbert projective metric, also referred to as the tropical metric, is given by
@@
The tropical projective torus is the ambient space containing the tropical Grassmannian, which is equivalent to the space of phylogenetic trees (Speyer & Sturmfels, (2004)). This has sparked the study of tropical data science, formalising statistical techniques which respect the tropical geometry of the state space (Monod et al., (2018); Yoshida, (2020)).
The metric above is widely accepted and preferred for the development of geometric statistical tools on the tropical projective torus (Yoshida, (2020)). However, recent work has shown theoretical advantages to the use of an asymmetric metric (Comăneci & Joswig, (2023)).
@@
The tropical asymmetric distance on $\mathbb{R}^{n}/\mathbb{R}\mathbf{1}$ is given by
@@
The asymmetric distances act as an $\ell_{1}$ norm on the tropical projective torus, which the symmetric metric $d_{\mathrm{tr}}$ acts as an $\ell_{\infty}$ norm. We note that $nd_{{\rm tr}}=d_{\triangle_{\min}}+d_{\triangle_{\max}}$. Figure 1 shows the balls for each tropical metric of interest.
@@
Definition 5 (Tropical Hyperplanes).
The tropical hyperplane $\mathcal{H}_{\mathbf{a}}$ defined by the linear form $a_{1}\odot x_{1}\boxplus\dots\boxplus a_{n}\odot x_{n}$ is given by
@@
Definition 6 (Tropical Line Segments).
@@
$A$
$B$$A$
$A$$B$$C$
@@
Definition 7 (Tropical Convex Hulls (Develin & Sturmfels,, 2004)).
A set $S$ is $\max$-tropically convex if it contains $\alpha\odot\mathbf{x}\boxplus\beta\odot\mathbf{y}$ for all $\mathbf{x},\mathbf{y}\in S$ and $\alpha,\beta\in\mathbb{R}$. For a finite subset $X=\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\}\subset\mathbb{R}^{n}/\mathbb{R}\mathbf{1}$, the max-plus tropical convex hull of $X$ is the set of all max-tropical linear combinations of points in $X$,
$\operatorname{tconv}_{\max}(X)=\{\alpha_{1}\odot\mathbf{x}_{1}\boxplus\alpha_{2}\odot\mathbf{x}_{2}\boxplus\cdots\boxplus\alpha_{n}\odot\mathbf{x}_{n}\mid\alpha_{1},\ldots,\alpha_{n}\in\mathbb{R}\}.$
Similarly, the min-tropical convex hull of $X$, $\operatorname{tconv}_{\min}(X)$ is the set of all min-tropical linear combinations of points in $X$.
A study of tropically convex functions requires an understanding of both max-tropical and min-tropical convex hulls, as we will see in our following consideration of tropical location problems.
@@ 2.2 Tropical Location Problems
Tropical location problems are a new avenue of research, motivated by the optimisation of statistical loss functions; they involve the minimisation over some dataset of some loss function which heuristically increases with distance from some kernel. Here we outline the relevant definitions of tropical location problems and their theoretical behaviour as presented by Comăneci, (2023). We introduce the problem of tropical linear regression (Akian et al., (2023)) as motivation.
@@ 2.2.1 Motivating Example
The problem of tropical linear regression looks to find the best-fit tropical hyperplane which minimises the maximal distance to a set of data points. While this is an inherently statistical optimisation problem, Akian et al., (2023) prove it to be polynomial-time equivalent to solving mean payoff games.
A tropical hyperplane is uniquely defined by a cone point $\mathbf{t}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$, allowing us to formulate the best-fit tropical hyperplane as an optimisation problem over $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$.
Definition 8 (Tropical Linear Regression (Akian et al., (2023))).
@@
In Figure 3, we see a heatmap of this objective function over the tropical projective plane. We see immediately that the objective is not convex. The linear regression objective has sparse gradient almost everywhere, which produces valleys. These properties are considered unfavourable extreme cases for gradient methods, but are particularly common in tropical optimisation problems. While the tropical linear regression problem is not convex, it is quasi-convex along tropical line segments; that is, its sub-level sets are tropically convex.
@@
2.2.2 $\triangle$-Star Quasi-Convexity
In formalising tropical location problems, Comăneci, (2023) first considers functions which increase along tropical geodesics from a kernel $\mathbf{v}$; that is, for any point $\mathbf{t}$ on a geodesic from $\mathbf{v}$ to $\mathbf{w}$, we have that $f(\mathbf{t})\leq f(\mathbf{w})$. This property is equivalent to star-convexity of sublevel sets. To formally state this, we define (oriented) geodesic segments and $\triangle_{\min}$-star convexity of sets.
@@ 2.2.2 $\triangle$-Star Quasi-Convexity
The (oriented) geodesic segment between $\mathbf{a},\mathbf{b}\in\mathbb{R}^{n}/\mathbb{R}\mathbf{1}$, is given by $[\mathbf{a},\mathbf{b}]_{\triangle_{\max}}\coloneqq\{\mathbf{x}\in\mathbb{R}^{n}/\mathbb{R}\mathbf{1}:d_{\triangle_{\max}}(\mathbf{a},\mathbf{x})+d_{\triangle_{\max}}(\mathbf{x},\mathbf{b})=d_{\triangle_{\max}}(\mathbf{a},\mathbf{b})\}$.
@@
A set $K\subseteq\mathbb{R}^{n}/\mathbb{R}\mathbf{1}$ is a $\triangle_{\max}$-star-convex set with kernel $\mathbf{v}$ if, for every $\mathbf{w}\in K$, we have $[\mathbf{v},\mathbf{w}]_{\triangle_{\max}}\subseteq K$.
Using the above definitions, we can characterise a loss function which is increasing along tropical geodesics as a function with $\triangle_{\max}$-star-convex sublevel sets. The following definition gives a closed form expression for such functions, while Theorem 12 proves their equivalence.
Definition 11 (Comăneci, (2023)).
A function $f:\mathbb{R}^{n}/\mathbb{R}\mathbf{1}\rightarrow\mathbb{R}$ is $\triangle_{\max}$-star-quasi-convex with kernel $\mathbf{v}$ if $f(\mathbf{x})=\hat{\gamma}(\mathbf{x}-\mathbf{v})$ for some increasing $\gamma:\mathbb{R}^{n}_{\geq 0}\rightarrow\mathbb{R}$, where $\hat{\gamma}(\mathbf{x})=\gamma((\max_{i}x_{i})\mathbf{1}-\mathbf{x})$.
Theorem 12 (Theorem 17, Comăneci, (2023)).
A continuous function $f:\mathbb{R}^{n}/\mathbb{R}\mathbf{1}\rightarrow\mathbb{R}$ is $\triangle_{\max}$-star-quasi-convex with kernel $\mathbf{v}$ if and only if all of its non-empty sub-level sets are $\triangle_{\max}$-star convex with kernel $\mathbf{v}$.
These $\triangle_{\max}$-star-convex functions act as a loss function with respect to a single datapoint. For example, the tropical metric $d_{\mathrm{tr}}$ is both $\triangle_{\min}$ and $\triangle_{\max}$-star-quasi-convex. The distance to a hyperplane $d_{\mathrm{tr}}(\mathbf{x},\mathcal{H}_{\mathbf{t}})$ is $\triangle_{\max}$-star-quasi-convex in $\mathbf{x}$, but $\triangle_{\min}$-star-quasi-convex in $\mathbf{t}$.
As a final remark, we note that $\triangle_{\max}$-star-quasi-convex functions are a special case of tropically quasi-convex functions by the following lemma. We limit our theoretical considerations to $\triangle_{\max}$-star-quasi-convex functions to utilise their closed form expression given in Definition 11, however numerical experiments demonstrate strong performance for more general tropical quasi-convex functions.
Lemma 13 (Comăneci, (2023)).
Any $\triangle_{\max}$-star-convex set is $\max$-tropically convex. Hence, any $\triangle_{\max}$-star-quasi-convex function is tropically quasiconvex in that it has tropically convex sublevel sets.

### v2

@@ Abstract
We propose a gradient descent method for solving optimization problems arising in settings of tropical geometry—a variant of algebraic geometry that has become increasingly studied in applications such as computational biology, economics, and computer science. Our approach takes advantage of the polyhedral and combinatorial structures arising in tropical geometry to propose a versatile approach for approximating local minima in tropical statistical optimization problems—a rapidly growing body of work in recent years. Theoretical results establish global solvability for 1-sample problems and a convergence rate matching classical gradient descent. Numerical experiments demonstrate the method’s superior performance over classical descent for tropical optimization problems which exhibit tropical convexity but not classical convexity. We also demonstrate the seamless integration of tropical descent into advanced optimization methods, such as Adam, offering improved overall performance.
@@ 1 Introduction
Where algebraic geometry studies geometric properties of solution sets of systems of multivariate polynomials, tropical geometry restricts to the case of polynomials defined by linearizing operations, where the “sum” of two elements is their maximum and the “product” of two elements is their sum. Evaluating polynomials with these operations results in piecewise linear functions. Tropical geometry is a relatively young field of pure mathematics established in the 1990s (with roots dating back further) and has recently become an area of active interest in computational and applied mathematics for its relevance to the statistical analyses of phylogenetic trees (Monod et al.,, 2018), the problem of dynamic programming in computer science (Maclagan & Sturmfels,, 2015), and in mechanism design for two-player games in game theory (Lin & Tran,, 2019). Following the identification of the tropical Grassmannian and the space of phylogenetic trees (Speyer & Sturmfels,, 2004) in particular, there has been surge of research (Barnhill et al.,, 2024b; Lin & Yoshida,, 2018; Comăneci & Joswig,, 2024; Monod et al.,, 2018) which necessitates the development of new computational techniques for optimisation in tropical geometric settings.
The training of machine learning models and the vast majority of computational tasks in statistics entail solving optimization problems; this paper is motivated by statistical optimization problems in tropical geometric settings. Initial research into tropical statistics considered optimization problems such as the identification of Fermat–Weber points which, due to the piecewise linear nature of tropical geometry, can be re-framed as a linear program (Comăneci & Joswig,, 2024; Lin & Yoshida,, 2018). Unfortunately, the complexity of these linear programs often scales with the size of our dataset, so gradient descent is more commonly implemented in practice (Barnhill et al.,, 2024a, b). More recently—following the identification of ReLU neural networks and tropical rational functions (Zhang et al.,, 2018)—more sophisticated neural network optimization problems are being studied in the tropical setting (Pasque et al.,, 2024; Yoshida et al.,, 2024), which necessitates the use of gradient methods for optimization in the tropical setting. In this paper, we address the natural question—given the need for gradient methods in tropical data science, how can we tailor classical gradient methods to the tropical setting? We provide a foundational tropical gradient descent framework to solve the optimization problems which arise in the computation of a wide range of tropical statistics, an area of active interest in applied tropical geometry.
@@
The remainder of this paper is organized as follows. The following section contains the necessary foundations of tropical geometry, an introduction to tropical location problems (Comăneci,, 2024), and the array of statistical optimization tasks studied in this paper. Section 3 contains the theoretical contributions of this paper; we formulate steepest descent with respect to the tropical norm and present the theoretical guarantees for tropical descent, most notably its convergence for a wide class of tropically quasi-convex functions and a convergence rate of $O(1/\sqrt{k})$. In Section 4, we perform a comprehensive numerical study of our proposed tropical descent method and its performance in the statistical optimization problems of interest. We conclude with a discussion of possible future work on tropical gradient methods in Section 5.
@@ 2 Background and Preliminaries
In this section, we overview the background on tropical geometry and present the statistical optimization problems which we study in the tropical setting.
@@ 2.1 Tropical Geometry
Here, we review the algebraic and geometric structure of the tropical projective torus, the state space for phylogenetic data and the domain of our tropical statistical optimization problems.
@@
The $n$-dimensional tropical projective torus is a quotient space constructed by endowing $\mathbb{R}^{N}$ with the equivalence relation
@@
it is denoted by $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$. The generalized Hilbert projective metric, also referred to as the tropical metric, is given by
@@
Remark 3.
While a single point in the tropical projective torus is given by some ray $\mathbf{x}+\mathbb{R}\mathbf{1}$, we note that there is a unique representative of this equivalence class whose coordinates sum to zero. In taking such representatives, we can identify the tropical projective torus $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ and the hyperplane =$\mathcal{H}=\{\mathbf{x}\in\mathbb{R}^{N}:\sum x_{i}=0\}$. Throughout this paper, we use this homeomorphism to visualize the tropical projective torus (e.g., Figures 1 and 2).
The tropical projective torus is the ambient space containing the tropical Grassmannian, which is equivalent to the space of phylogenetic trees (Speyer & Sturmfels,, 2004). This has sparked the study of tropical data science, formalizing statistical techniques which respect the tropical geometry of the state space (Monod et al.,, 2018; Yoshida,, 2022).
The generalized Hilbert projective metric defined above is widely accepted and preferred for the development of geometric statistical tools on the tropical projective torus (Yoshida,, 2022). However, recent work has shown theoretical advantages to the use of an asymmetric metric (Comăneci & Joswig,, 2024).
@@
The tropical asymmetric distances on $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ are given by
@@
The asymmetric distances act as an $\ell_{1}$ norm on the tropical projective torus, while the symmetric metric $d_{\mathrm{tr}}$ acts as an $\ell_{\infty}$ norm. We note that $Nd_{{\rm tr}}=d_{\triangle_{\min}}+d_{\triangle_{\max}}$. Figure 1 shows the balls for each tropical metric of interest.
@@
Definition 6 (Tropical Hyperplane).
The tropical hyperplane $\mathcal{H}_{\mathbf{a}}$ defined by the linear form $a_{1}\odot x_{1}\boxplus\cdots\boxplus a_{N}\odot x_{N}$ is given by
@@
Definition 7 (Tropical Line Segment).
@@
$\mathbf{a}$
$\mathbf{b}$$\mathbf{a}$
$\mathbf{a}$$\mathbf{b}$$\mathbf{c}$
@@
Definition 8 (Tropical Convex Hull (Develin & Sturmfels,, 2004)).
A set $S\subset\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ is $\max$-tropically convex if it contains $\alpha\odot\mathbf{x}\boxplus\beta\odot\mathbf{y}$ for all $\mathbf{x},\mathbf{y}\in S$ and $\alpha,\beta\in\mathbb{R}$. For a finite subset $X=\{\mathbf{x}_{1},\ldots,\mathbf{x}_{K}\}\subset\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$, the max-plus tropical convex hull of $X$ is the set of all max-tropical linear combinations of points in $X$,
$\operatorname{tconv}_{\max}(X)\coloneqq\{\alpha_{1}\odot\mathbf{x}_{1}\boxplus\cdots\boxplus\alpha_{K}\odot\mathbf{x}_{K}\mid\alpha_{1},\ldots,\alpha_{K}\in\mathbb{R}\}.$
Similarly, the min-tropical convex hull of $X$, $\operatorname{tconv}_{\min}(X)$, is the set of all min-tropical linear combinations of points in $X$.
A study of tropically convex functions requires an understanding of both max-tropical and min-tropical convexity, as we will see in our following consideration of tropical location problems.
@@ 2.2 Tropical Location Problems
Tropical location problems are a new avenue of research, motivated by the optimization of statistical loss functions; they involve the minimization over some dataset of some loss function which heuristically increases with distance from some kernel. Here we outline the relevant definitions of tropical location problems and their theoretical behavior as presented by Comăneci, (2024). We introduce the problem of tropical linear regression (Akian et al.,, 2023) as motivation.
@@ 2.2.1 Motivating Example
The problem of tropical linear regression looks to find the best-fit tropical hyperplane which minimizes the maximal distance to a set of data points. While this is an inherently statistical optimization problem, it has recently been proven to be polynomial-time equivalent to solving mean payoff games (Akian et al.,, 2023).
A tropical hyperplane is uniquely defined by a cone point $\mathbf{t}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$, allowing us to formulate the best-fit tropical hyperplane as an optimization problem over $\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$.
Definition 9 (Tropical Linear Regression (Akian et al.,, 2023)).
@@
In Figure 3, we see a heat map of this objective function over the tropical projective plane for a sample of size 3. We see immediately that the objective is not convex. The linear regression objective has sparse gradient almost everywhere, which produces valleys. These properties are considered unfavourable extreme cases for gradient methods, but are particularly common in tropical optimization problems. While the tropical linear regression problem is not convex, it is quasi-convex along tropical line segments; that is, its sub-level sets are tropically convex.
@@
2.2.2 Quasi-Convexity of Tropical Functions
In formalizing tropical location problems, the work of Comăneci, (2024) first considers functions which increase along tropical geodesics from a kernel $\mathbf{v}$; that is, for any point $\mathbf{t}$ on a geodesic from $\mathbf{v}$ to $\mathbf{w}$, we have that $f(\mathbf{t})\leq f(\mathbf{w})$. This property is equivalent to star-convexity of sublevel sets. To formally state this, we define (oriented) geodesic segments and $\triangle_{\min}$-star convexity of sets.
@@ 2.2.2 Quasi-Convexity of Tropical Functions
The (oriented) geodesic segment between $\mathbf{a},\mathbf{b}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$, is given by $[\mathbf{a},\mathbf{b}]_{\triangle_{\min}}\coloneqq\{\mathbf{x}\in\mathbb{R}^{N}/\mathbb{R}\mathbf{1}:d_{\triangle_{\min}}(\mathbf{a},\mathbf{x})+d_{\triangle_{\min}}(\mathbf{x},\mathbf{b})=d_{\triangle_{\min}}(\mathbf{a},\mathbf{b})\}$.
The oriented geodesic segment is the union of all geodesics from $\mathbf{a}$ to $\mathbf{b}$ with respect to $d_{\triangle_{\min}}$.
@@
A set $K\subseteq\mathbb{R}^{N}/\mathbb{R}\mathbf{1}$ is a $\triangle_{\min}$-star-convex set with kernel $\mathbf{v}$ if, for every $\mathbf{w}\in K$, we have $[\mathbf{v},\mathbf{w}]_{\triangle_{\min}}\subseteq K$.
Using the above definitions, we can characterize a loss function which is increasing along tropical geodesics as a function with $\triangle_{\min}$-star-convex sublevel sets. The following definition gives a closed form expression for such functions, while Theorem 13 proves their equivalence.
Definition 12 ($\triangle_{\min}$-Star-Quasi-Convex Functions (Comăneci,, 2024)).
A function $f:\mathbb{R}^{N}/\mathbb{R}\mathbf{1}\rightarrow\mathbb{R}$ is $\triangle_{\min}$-star-quasi-convex with kernel $\mathbf{v}$ if $f(\mathbf{x})=\hat{\gamma}(\mathbf{x}-\mathbf{v})$ for some $\gamma:\mathbb{R}^{N}_{\geq 0}\rightarrow\mathbb{R}$ which is increasing in every coordinate, where we define $\hat{\gamma}(\mathbf{x})\coloneqq\gamma(\mathbf{x}-(\min_{i}x_{i})\mathbf{1})$.
Theorem 13 (Theorem 17 of (Comăneci,, 2024)).
A continuous function $f:\mathbb{R}^{N}/\mathbb{R}\mathbf{1}\rightarrow\mathbb{R}$ is $\triangle_{\min}$-star-quasi-convex with kernel $\mathbf{v}$ if and only if all of its non-empty sub-level sets are $\triangle_{\min}$-star convex with kernel $\mathbf{v}$.
Remark 14.
We can define $\triangle_{\max}$-star-quasi-convex functions similarly, but we also note that a function $f(\mathbf{t})$ is a $\triangle_{\max}$-star-quasi-convex function if and only if $f(-\mathbf{t})$ is a $\triangle_{\min}$-star-quasi-convex function.
These $\triangle_{\min}$-star-convex functions act as a loss function with respect to a single data point. For example, the tropical metric $d_{\mathrm{tr}}$ is both $\triangle_{\min}$ and $\triangle_{\max}$-star-quasi-convex. The distance to a hyperplane $d_{\mathrm{tr}}(\mathbf{x},\mathcal{H}_{\mathbf{t}})$ is $\triangle_{\max}$-star-quasi-convex in $\mathbf{x}$, but $\triangle_{\min}$-star-quasi-convex in $\mathbf{t}$.
As a final remark, we note that $\triangle_{\min}$-star-quasi-convex functions are a special case of tropically quasi-convex functions by the following lemma. We limit our theoretical considerations to $\triangle_{\min}$-star-quasi-convex functions to utilize their closed form expression given in Definition 12, however numerical experiments demonstrate strong performance of our methodology for more general tropical quasi-convex functions.
Proposition 15 (Proposition 8 of (Comăneci,, 2024)).
Any $\triangle_{\min}$-star-convex set is $\min$-tropically convex. Hence, any $\triangle_{\min}$-star-quasi-convex function is tropically quasi-convex in that it has tropically convex sublevel sets.